OpenStudy (annon):

a particle in the xy-plane starts at the point (4,2) and moves along the curve y= sqrt(x) to the point (0,0) and then returns along the same curve to point(1,1). use a single vector-valued function to represent the path of the particle over a single interval (for the parameter). dont forget to give the interval for the parameter.

OpenStudy (anonymous):

x=t^2 / 4 y=t/2

OpenStudy (annon):

interval?

OpenStudy (anonymous):

t goes frm 4 to 0 on first trip from 0 to 1 on the second

OpenStudy (annon):

please :)

OpenStudy (annon):

isnt the (t) function suppose to be continuious?

OpenStudy (anonymous):

isnt this continuous?

OpenStudy (dumbcow):

i think it should be 0 to 2 for 2nd trip: 2^2/4 = 1, 2/2 = 1

OpenStudy (anonymous):

oh ys..my bad

OpenStudy (dumbcow):

and i agree it is continuous

OpenStudy (annon):

im sorry i ment i thought that there should be one interval for both equations